OpenStudy (anonymous):

What is the inverse function of f(x) = 4 - x/6 ?

OpenStudy (anonymous):

x=4-(f(x))/6 (f(x))/6=-x+4 f(x)=-6x+24

OpenStudy (anonymous):

thank you!!! :)

OpenStudy (anonymous):

is it \[f(x)=4-\frac{x}{6}\] or \[f(x)=\frac{4-x}{6}\]?